Although the Cheetah is, without a doubt, the fastest ground runner, it can only maintain an incredible speed of 110 km/h for about 175 meters.
a) how long is that?
b) If a cheetah is in chase after a gazelle running at 50 km/h, how far should it be from its prey to guarantee a fruitful chase? Assume both maintain constant speeds during chase.
1
Expert's answer
2019-10-07T11:04:43-0400

a) convert the speed into SI units:

"110 km/h = 110\cdot 1000/3600\, m/s = 30.56\, m/s"

as we know

"t=\frac{S}{v} = \frac{175}{30.56} = 5.73\, s"

b) "S=(v_c-v_g)\cdot t = (110-50)\cdot5.73/3.6 = 95.5\, m"